jwbejflm:

Given the information below, write the equation of each line in slope intercept form. Slope - 1 /4 and going through the point (2, 6).

darkknight:

the slope is 1/4? so \[y - 6 = 1/4(x-2)\] so then y -6 = 1/4(x) - 2/4 Then y = 1/4(x) + 22/4

jwbejflm:

do you need the picture of the graph

darkknight:

does the slope say 1/4 or -1/4?

jwbejflm:

Slope - 1 /4 and going through the point (2, 6).

darkknight:

is that a negative slope or a positive??? Can you post the graph?

jwbejflm:

darkknight:

alright, so the slope was negaitve. So first use point slope form \[y - 6 = -1/4(x-2)\] Can you convert to slope-intercept from here?

jwbejflm:

not really

darkknight:

okay so first distribute in the parenthesis, then add 6 to both sides to solve for y

jwbejflm:

so y−12=−1/4x−8 ?

darkknight:

y−6=−1/4(x−2) so y - 6 = -1/4x + 2/4 Now -1/4x + 26/4 = y

jwbejflm:

okay thank you